OpenStudy (anonymous):

Solve the following system using the substitution method. ----->>>> 3x+y=-3 y=2x+7 coordinate (___,___) ??

OpenStudy (anonymous):

(-2,3) ??

OpenStudy (anonymous):

we know that y = 2x + 7 so sub 2x + 7 in for y in the other equation 3x + y = -3 3x + 2x + 7 = -3 5x + 7 = -3 5x = -3 - 7 5x = -10 x = -2 now sub -2 in for x in either of the equations y = 2x + 7 y = 2(-2) + 7 y = -4 + 7 y = 3 check... 3x + y = -3 3(-2) + 3 = -3 -6 + 3 = -3 -3 = -3 (correct) x = -2 and y = 3 (-2,3) any questions ?

OpenStudy (anonymous):

yes (-2,3)

OpenStudy (anonymous):

thank you